# Non-Rigid Designators in Modal and Temporal Free Description Logics

### v1

@@ Abstract
Definite descriptions, such as ‘the General Chair of KR 2024’, are a semantically transparent device for object identification in knowledge representation. In first-order modal logic, definite descriptions have been widely investigated for their non-rigidity, which allows them to designate different objects (or none at all) at different states. We propose expressive modal description logics with non-rigid definite descriptions and names, and investigate decidability and complexity of the satisfaction problem. We first systematically link satisfiability for the one-variable fragment of first-order modal logic with counting to our modal description logics. Then, we prove a promising NExpTime-completeness result for concept satisfiability for the fundamental epistemic multi-agent logic $\mathbf{S5}^{n}$ and its neighbours, and show that some expressive logics that are undecidable with constant domain become decidable (but Ackermann-hard) with expanding domains. Finally, we conduct a fine-grained analysis of decidability of temporal logics.
@@ 1 Introduction
Definite descriptions, like ‘the General Chair of KR 2024’, are expressions of the form ‘the $x$ such that $\varphi$.’ Together with individual names such as ‘Pierre,’ they are used as referring expressions to identify objects in a given domain (???).
@@
In particular, DLs $\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}$ and $\smash{\mathcal{ELO}_{\!u}^{\smash{\iota}}}$ (??) include the universal role, $u$, as well as nominals and definite descriptions of the form $\{\smash{\iota}C\}$ (‘the object in $C$’) as basic concept constructs, while also employing a free logic semantics that allows non-designating terms. Hence, for instance, using nominal $\{\mathsf{kr24}\}$ to designate KR 2024, we can refer to the General Chair of KR 2024 by means of the definite description $\{\smash{\iota}\exists\mathsf{isGenChair}.\{\mathsf{kr24}\}\}$. Then, in $\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}$, we can say that Pierre is the General Chair of KR 2024 with the following concept:
@@
The free DLs $\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}$ and $\smash{\mathcal{ELO}_{\!u}^{\smash{\iota}}}$ have, respectively, ExpTime-complete ontology satisfiability and PTime-complete entailment problems, thus matching the complexity of the classical DL counterparts.
@@
The aim of this contribution is to introduce modal DLs that have sufficient expressive power to represent the phenomena discussed above, explore their relationship to standard modal DLs without definite descriptions and non-designating names, and investigate decidability and complexity of reasoning in these languages.
@@
In a temporal setting, where $\Box$ is read as ‘always’ and its dual $\Diamond$ as ‘at some point in the future’, KR 2024 can be made a rigid designator, which refers to the same object at all time instants. This can be achieved, for example, by means of an ontology that holds globally, at all time instants, and consists of the following concept inclusion (CI):
@@
Compared to first-order modal logics with non-rigid designators (??), both the definition of the language and the scope distinctions for modal operators are simplified in modalised DLs, as first-order variables are replaced by a class-based DL syntax that leaves quantification and predicate abstraction implicit. It is, however, possible to translate our modal DL back to a natural fragment of first-order modal logic with definite descriptions and predicate abstraction.
@@
In this work, we consider interpretations with both constant domains (in which the first-order domain is fixed across all worlds) and expanding domains (in which the domain can grow when moving to accessible worlds). We first establish, for any class of Kripke frames and both constant and expanding domains, polytime satisfiability-preserving reductions (with and without ontology) to the language $\smash{\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}}$ without definite descriptions. We will show that, in addition, we can assume that each nominal designates in every world, but importantly is still non-rigid.
@@
We then study the satisfiability problem for various fundamental modal logics with epistemic and temporal interpretations. While for first-order modal logics with only rigid designators and no counting the restriction to monodic formulas (in which modal operators are only applied to formulas with a single free variable) very often ensures decidability, this is no longer the case if non-rigid designators and/or some counting are admitted (?). For our modal DLs, this implies that the standard recipe for designing decidable languages — apply modal operators only to concepts — does not always work anymore. Here, we explore in detail when this recipe still works, and when it does not.
@@
First, we closely link the two main sources of bad computational behaviour, non-rigid designators and counting, enabling us to use the results and machinery introduced for first-order modal logics with counting (???). Then, we prove that, rather surprisingly, for some fundamental modal epistemic logic, non-rigid designators come for free: concept satisfiability for the modal logics of all Kripke frames with $n$ accessibility relation, $\mathbf{K}^{n}$, and of all Kripke frames with $n$ equivalence relations, $\mathbf{S5}^{n}$, is in NExpTime and thus no harder than without names at all. This holds under both constant and expanding domains, and the proof is by showing the exponential finite model property. This answers an open problem discussed in (?). With ontologies, however, concept satisfiability becomes undecidable under constant domains. While for $\mathbf{S5}^{n}$ (because of symmetric accessibility relations) constant domains coincide with expanding domains, for $\mathbf{K}^{n}$ decidability under expanding domains remains open with ontologies. As a fundamental example of an expressive modal logic, we investigate the extension $\mathbf{K}^{\ast n}$ of $\mathbf{K}^{n}$ with a modal operator interpreted by the transitive closure of the union of the $n$ accessibility relations, which can be interpreted as common knowledge (?) but also as a fragment of propositional dynamic logic, PDL (?). In this case concept satisfiability is undecidable under expanding and constant domains, but becomes decidable, though Ackermann-hard, for the corresponding logic, $\mathbf{K}\!\boldsymbol{f}^{\ast n}$, on finite acyclic models with expanding domains. This answers an open problem posed in (?). Note that Ackermann-hardness means that the time required to establish (un)satisfiability is not bounded by any primitive recursive (or computable) function. We refer the reader to Table 1 for an overview of our results. Recall that with rigid designators all these problems are known to be in 2NExpTime (??).
@@
$\mathbf{K}^{n}$, $n\geq 1$ const. NExp-c. undecidable
exp. NExp-c. ?
$\mathbf{S5}$ NExp-complete
$\mathbf{S5}^{n}$, $n\geq 2$ NExp-c. undecidable
$\mathbf{K}^{\ast n}$, $n\geq 1$ const. $\Sigma^{1}_{1}$-complete
exp. undecidable
$\mathbf{K}\!\boldsymbol{f}^{\ast n}$, $n\geq 1$ const. undecidable
exp. decidable, Ackermann-hard
@@
It is to be emphasised that the non-rigidity of symbol interpretation by itself is not the cause for the satisfiability problem to become harder. For instance, rigid roles are known to often cause harder satisfaction problems than non-rigid roles (?). What makes non-rigid designators computationally much harder than rigid designators is their ability to count in an unbounded way across worlds.
@@
Other than in non-modal DL languages (???), and in the already mentioned first-order modal, temporal and hybrid logic settings, definite descriptions have been recently investigated in the context of propositional modal, temporal and hybrid logics with nominals and the @@@@ operator (?). Here, the additional $\smash{\iota}$ operator allows one to refer to the (one and only) state of a model that satisfies a certain condition.
Non-rigid rigid designators have received, to the best of our knowledge, little attention in modal DLs, despite the wide body of research both on temporal (????) and epistemic (????) extensions. As a notable exception, in an epistemic DL context, non-rigid individual names appear in (?), where abstract individual names are interpreted on an infinite common domain, but without definite descriptions.
@@ 2 Preliminaries
A frame is a pair $\mathfrak{F}=(W,\{R_{i}\}_{i\in I})$, where $W$ is a non-empty set of worlds (or states) and each $R_{i}\subseteq W\times W$, for $i\in I$, is a binary accessibility relation on $W$. A partial interpretation with expanding domains based on a frame $\mathfrak{F}=(W,\{R_{i}\}_{i\in I})$ is a triple $\mathfrak{M}=(\mathfrak{F},\Delta,\mathcal{I})$, where $\Delta$ is a function associating with every $w\in W$ a non-empty set, $\Delta^{w}$, called the domain of $w$ in $\mathfrak{M}$, such that $\Delta^{w}\subseteq\Delta^{v}$, whenever $wR_{i}v$, for some $i\in I$; and $\mathcal{I}$ is a function associating with every $w\in W$ a partial DL interpretation $\mathcal{I}_{w}=(\Delta^{w},\cdot^{\mathcal{I}_{w}})$ that maps every $A\in{\sf N_{C}}$ to a subset $A^{\mathcal{I}_{w}}$ of $\Delta^{w}$, every $r\in{\sf N_{R}}$ to a subset $r^{\mathcal{I}_{w}}$ of $\Delta^{w}\times\Delta^{w}$, the universal role $u$ to the set $\Delta^{w}\times\Delta^{w}$, and every $a$ in some subset of ${\sf N_{I}}$ to an element $a^{{\mathcal{I}_{w}}}$ in $\Delta^{w}$. In particular, every $\cdot^{\mathcal{I}_{w}}$ is a total function on ${\sf N_{C}}\cup{\sf N_{R}}$ but a partial function on ${\sf N_{I}}$. If $\mathcal{I}_{w}$ is defined on $a\in{\sf N_{I}}$, then we say that $a$ designates at $w$. If every $a\in{\sf N_{I}}$ designates at $w\in W$, then $\mathcal{I}_{w}$ is called total. We say that $\mathfrak{M}=(\mathfrak{F},\Delta,\mathcal{I})$ is a total interpretation if every $\mathcal{I}_{w}$, $w\in W$, is a total interpretation. In the sequel, we refer to partial interpretations as interpretations, and add the adjective ‘total’ explicitly whenever this is the case.
@@
To avoid ambiguities, we use square brackets when applying negation and modal operators to assertions, as in $\lnot[C(\tau)]$ and $\Diamond_{i}[C(\tau)]$. Observe that, in an assertion of the form $\Diamond_{i}C(\tau)$, the diamond acts as a de re operator, since the concept $\Diamond_{i}C$ applies to the object, if any, designated by the term $\tau$ at the current world $w$, and the assertion is false at a world whenever $\tau$ fails to designate at $w$. On the other hand, in an expression of the form $\Diamond_{i}[C(\tau)]$, the diamond plays the role of a de dicto modality, as it refers to the world of evaluation for the whole assertion $C(\tau)$. Using the lambda abstraction notation for first-order modal logic (?), assertion $\Diamond_{i}A(a)$ corresponds to $\exists x.({\langle}\lambda y.x=y{\rangle}(a)\land\Diamond_{i}A(x))$, whereas $\Diamond_{i}[A(a)]$ stands for $\Diamond_{i}\exists x.({\langle}\lambda y.x=y{\rangle}(a)\land A(x))$; see Appendix A for details.
@@ 3 Reasoning Problems and Reductions
In the sequel, for the case of concept $\mathcal{C}$-satisfiability under global ontology, we will assume without loss of generality that $C$ is a concept name. Indeed, we can extend $\mathcal{O}$ with $A\equiv C$, for a fresh concept name $A$, and consider satisfiability of $A$ under the extended ontology, which is a model conservative extension of $\mathcal{O}$.
We begin with a few observations on polytime reductions between the concept satisfiability problems (under global ontology) for different languages and semantic conditions, including (non-)rigid designators, total and partial interpretations, definite description and nominals, the universal role, and expanding and constant domains. These observations will be useful in our constructions below. All proofs are available in the appendix.
@@
is unsatisfiable in interpretations with the RDA, but is satisfiable otherwise as $a$ can designate differently at different worlds. Note that an individual that fails to designate at all worlds is vacuously rigid. The following proposition shows the RDA can be enforced in interpretations by an ontology.
@@
The proof is based on the observation that CIs of the form $\{a\}\sqsubseteq\Box_{i}\{a\}$, for $i\in I$, ensure that $a$ can be made a rigid designator in any interpretation. This provides a reduction for the case of global ontology, where a given $\mathcal{O}$ is extended with these CIs for every $a$. A similar reduction is provided in Appendix B for concept satisfiability in total interpretations.
@@
We sketch the proof for the case of global ontology. Let $\mathcal{O}$ be an ontology. Consider the extension $\mathcal{O}^{\prime}$ of $\mathcal{O}$ with CIs $\top\sqsubseteq\exists u.\{a\}$, for all individual names $a$ in $\mathcal{O}$. Clearly, these CIs ensure that each $a$ designates in every accessible world. The case of concept satisfiability is shown in the appendix.
@@
We again sketch the proof for the case of global ontology. Let $\mathcal{O}$ be an ontology. Consider $\mathcal{O}^{\prime}$ obtained from $\mathcal{O}$ by replacing every nominal $\{a\}$ in $\mathcal{O}$ with a fresh concept name $N_{a}$ and by extending the result with all CIs $N_{a}\sqsubseteq\{a\}$. It follows that every $a$ in $\mathcal{O}^{\prime}$ can designate in all worlds, but the corresponding concept $N_{a}$ may still be interpreted by the empty set in some worlds, thus reflecting the fact that $a$ in $\mathcal{O}$ could have failed to designate in those worlds. The case of concept satisfiability is dealt with in the appendix.
@@
Next, we define normal form that will help us prove further polytime-reductions, e.g., Lemma 6 and Proposition 8, and then complexity upper bounds. Let $\mathcal{O}$ be an ontology and $C$ a subconcept in $\mathcal{O}$. Denote by $\mathcal{O}[C/A]$ the result of replacing every occurrence of $C$ in $\mathcal{O}$ with a fresh concept name $A$, called the surrogate of $C$. Clearly, $\mathcal{O}[C/A]\cup\{C\equiv A\}$ is a model conservative extension of $\mathcal{O}$. We can systematically apply this procedure to obtain an ontology in normal form where connectives are applied only to concept names: e.g., definite descriptions occur only in the form of $\smash{\iota}B$, for a concept name $B$. If surrogates are introduced for innermost connectives first, then the transformation runs in polytime.
@@
We first observe that nominals can be easily encoded with definite descriptions. For instance, given an ontology $\mathcal{O}$, take a fresh concept name $N_{a}$ for each individual name $a$ in $\mathcal{O}$, and let $\mathcal{O}^{\prime}$ be the result of replacing every occurrence of $\{a\}$ in $\mathcal{O}$ with $\{\smash{\iota}N_{a}\}$. Clearly, $\mathcal{O}^{\prime}$ is a model conservative extension of $\mathcal{O}$, and vice versa. Note, however, that $\mathcal{O}^{\prime}$ contains no nominals, and so there is neither a distinction between partial and total interpretations, nor between the RDA and no-RDA cases. In general, we have the following result.
@@
where $a_{B}$ is a fresh nominal. Intuitively, the first CI ensures that the surrogate for $\{\smash{\iota}B\}$ belongs to $B$ and is never interpreted by more than one element. The second CI ensures that if $B$ is a singleton, then that element belongs to the surrogate for $\{\smash{\iota}B\}$. Formally, we show that $\mathcal{O}^{\ast}$ is a model conservative extension of $\mathcal{O}$. The case of concept satisfiability relies on normal form of concepts (Lemma 5) and is treated in the appendix.
@@
It is known that, for the satisfiability problems, the interpretations with expanding domains can be simulated by constant domain interpretations using a fresh concept name representing the domain to relativise concepts and CIs; see e.g., (?, Proposition 3.32 (ii), (iv)). We restate this standard result in our setting for completeness:
@@
In $\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}$, concept $\mathcal{C}$-satisfiability (under global ontology) with expanding domains is polytime-reducible to concept $\mathcal{C}$-satisfiability (under global ontology, respectively) with constant domain.

### v2

@@ Abstract
Definite descriptions, such as ‘the General Chair of KR 2024,’ are a semantically transparent device for object identification in knowledge representation. In first-order modal logic, definite descriptions have been widely investigated for their non-rigidity, which allows them to designate different objects (or none at all) at different states. We propose expressive modal description logics with non-rigid definite descriptions and names, and investigate decidability and complexity of the satisfiability problem. We first systematically link satisfiability for the one-variable fragment of first-order modal logic with counting to our modal description logics. Then, we prove a promising NExpTime-completeness result for concept satisfiability for the fundamental epistemic multi-agent logic $\mathbf{S5}^{n}$ and its neighbours, and show that some expressive logics that are undecidable with constant domain become decidable (but Ackermann-hard) with expanding domains. Finally, we conduct a fine-grained analysis of decidability of temporal logics.
@@ 1 Introduction
Definite descriptions, like ‘the General Chair of KR 2024,’ are expressions of the form ‘the unique $x$ such that $\varphi$.’ Together with individual names such as ‘Pierre,’ they are used as referring expressions to identify objects in a given domain (???).
@@
In particular, DLs $\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}$ and $\smash{\mathcal{ELO}_{\!u}^{\smash{\iota}}}$ (??) include the universal role, $u$, as well as nominals and definite descriptions of the form $\{\smash{\iota}C\}$ (‘the unique object in $C$’) as basic concept constructs, while also employing a free logic semantics that allows non-designating terms. Hence, for instance, using nominal $\{\mathsf{kr24}\}$ to designate KR 2024, we can refer to the General Chair of KR 2024 by means of the definite description $\{\smash{\iota}\exists\mathsf{isGenChair}.\{\mathsf{kr24}\}\}$. Then, in $\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}$, we can say that Pierre is the General Chair of KR 2024 with the following concept:
@@
Indeed, if this concept is satisfiable (in other words, has a non-empty extension), then there are objects $p$ and $k$ designated by $\mathsf{pierre}$ and $\mathsf{kr24}$, respectively, and the pair $(p,k)$ belongs to the interpretation of role $\mathsf{isGenChair}$, which contains no other $(o,k)$, thus making $p$ the only object connected to $k$ by $\mathsf{isGenChair}$. The free DLs $\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}$ and $\smash{\mathcal{ELO}_{\!u}^{\smash{\iota}}}$ have, respectively, ExpTime-complete ontology satisfiability and PTime-complete entailment problems, thus matching the complexity of the classical DL counterparts.
@@
The aim of this contribution is to introduce modal DLs that have sufficient expressive power to represent the phenomena discussed above, explore their relationship to standard modal DLs without definite descriptions and non-designating names, and investigate decidability and complexity of reasoning in these formalisms.
@@
In a temporal setting, where $\Box$ is read as ‘always’ and its dual $\Diamond$ as ‘at some point in the future,’ KR 2024 can be made a rigid designator, which refers to the same object at all time instants. This can be achieved, for example, by means of an ontology that holds globally, at all time instants, and consists of the following concept inclusion (CI):
@@
Compared to first-order modal logics with non-rigid designators (??), both the definition of the language and the scope distinctions for modal operators are simplified in modalised DLs, as first-order variables are replaced by a class-based DL syntax leaving quantification and predicate abstraction implicit. We, however, show that our modal DLs can be translated to a natural fragment of first-order modal logic with definite descriptions and predicate abstraction.
@@
In this work, we consider interpretations with both constant domains (in which the first-order domain is fixed across all worlds) and expanding domains (in which it can grow when moving to accessible worlds). We first establish, for any class of Kripke frames and both constant and expanding domains, polytime satisfiability-preserving reductions (with and without ontology) to the language $\smash{\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}}$ without definite descriptions. We will show that, in addition, we can assume that each nominal designates in every world, but importantly is still non-rigid.
@@
We then study the satisfiability problem for various fundamental modal logics with epistemic and temporal interpretations. While for first-order modal logics with rigid designators and no counting the restriction to monodic formulas (in which modal operators are applied only to formulas with at most one free variable) very often ensures decidability, this is no longer the case if non-rigid designators and/or some counting are admitted (?). For our modal DLs, this implies that the standard recipe for designing decidable languages — apply modal operators only to concepts — does not always work anymore. Here, we explore in detail when this recipe still works, and when it does not.
First, we closely link the two main sources of bad computational behaviour, non-rigid designators and counting, enabling us to use the results and machinery introduced for first-order modal logics with counting (???).
@@
Next, we prove that, rather surprisingly, for some fundamental modal epistemic logic, non-rigid designators come for free: concept satisfiability for the modal logics of all Kripke frames with $n$ accessibility relation, $\mathbf{K}^{n}$, and of all Kripke frames with $n$ equivalence relations, $\mathbf{S5}^{n}$, is in NExpTime and thus no harder than without names at all. This holds under both constant and expanding domains, and the proof is by showing the exponential finite model property. This answers an open problem discussed in (?). With ontologies, however, concept satisfiability becomes undecidable under constant domains. While for $\mathbf{S5}^{n}$ (because of symmetric accessibility relations) constant domains coincide with expanding domains, for $\mathbf{K}^{n}$ decidability under expanding domains remains open with ontologies. As a fundamental example of an expressive modal logic, we investigate the extension $\mathbf{K}^{\ast n}$ of $\mathbf{K}^{n}$ with a modal operator interpreted by the transitive closure of the union of the $n$ accessibility relations, which can be interpreted as common knowledge (?) but also as a fragment of propositional dynamic logic, PDL (?). In this case concept satisfiability is undecidable under expanding and constant domains, but becomes decidable, though Ackermann-hard, for the corresponding logic, $\mathbf{K}\!\boldsymbol{f}^{\ast n}$, on finite acyclic models with expanding domains. This answers an open problem posed in (?). Note that Ackermann-hardness means that the time required to establish (un)satisfiability is not bounded by any primitive recursive (or computable) function. We refer the reader to Table 1 for an overview of our results. Recall that with rigid designators all these problems are known to be in 2NExpTime (??).
@@
$\mathbf{K}^{n}$, $n\geq 1$ const. NExp-c. [T 11] undecidable [T 14]
exp. NExp-c. [T 11] ?
$\mathbf{S5}$ NExp-complete [T 11]
$\mathbf{S5}^{n}$, $n\geq 2$ NExp-c. [T 11] undecidable [T 14]
$\mathbf{K}^{\ast n}$, $n\geq 1$ const. $\Sigma^{1}_{1}$-complete [L 13 + T 15]
exp. undecidable [L 13 + T 16.1]
$\mathbf{K}\!\boldsymbol{f}^{\ast n}$, $n\geq 1$ const. undecidable [L 13 + T 15]
exp.
decidable [T 12],
Ackermann-hard [L 13 + T 16.2]
@@
It is to be emphasised that the non-rigidity of symbol interpretation by itself is not the cause for the satisfiability problem to become harder. For instance, rigid roles are known to often cause an increase in the hardness of the satisfiability problem compared with the case of non-rigid roles only (?). What makes non-rigid designators computationally much harder than rigid designators is their ability to count in an unbounded way across worlds.
Full proofs and additional material are available in the appendix.
@@
Other than in non-modal DLs (???), and in the already mentioned first-order modal and temporal settings, definite descriptions have been recently investigated in the context of propositional hybrid logics with nominals and the @@@@ operator (?). Here, the additional $\smash{\iota}$ operator allows one to refer to the (one and only) state of a model that satisfies a certain condition.
Non-rigid designators have received, to the best of our knowledge, little attention in modal DLs, despite the extensive body of research both on temporal (????) and epistemic (????) extensions. As a notable exception, ? (?) investigate non-rigid individual names in an epistemic DL context, where abstract individual names are interpreted on an infinite common domain, but without definite descriptions.
@@ 2 Preliminaries
A frame is a pair $\mathfrak{F}=(W,\{R_{i}\}_{i\in I})$, where $W$ is a non-empty set of worlds (or states) and each $R_{i}\subseteq W\times W$, for $i\in I$, is a binary accessibility relation on $W$. A partial interpretation with expanding domains based on a frame $\mathfrak{F}=(W,\{R_{i}\}_{i\in I})$ is a triple $\mathfrak{M}=(\mathfrak{F},\Delta,\mathcal{I})$, where $\Delta$ is a function associating with every $w\in W$ a non-empty set, $\Delta^{w}$, called the domain of $w$ in $\mathfrak{M}$, such that $\Delta^{w}\subseteq\Delta^{v}$, whenever $wR_{i}v$, for some $i\in I$; and $\mathcal{I}$ is a function associating with every $w\in W$ a partial DL interpretation $\mathcal{I}_{w}=(\Delta^{w},\cdot^{\mathcal{I}_{w}})$ that maps every $A\in{\sf N_{C}}$ to a subset $A^{\mathcal{I}_{w}}$ of $\Delta^{w}$, every $r\in{\sf N_{R}}$ to a subset $r^{\mathcal{I}_{w}}$ of $\Delta^{w}\times\Delta^{w}$, the universal role $u$ to the set $\Delta^{w}\times\Delta^{w}$, and every $a$ in some subset of ${\sf N_{I}}$ to an element $a^{{\mathcal{I}_{w}}}$ in $\Delta^{w}$. Hence, every $\cdot^{\mathcal{I}_{w}}$ is a total function on ${\sf N_{C}}\cup{\sf N_{R}}$ but a partial function on ${\sf N_{I}}$. If $\mathcal{I}_{w}$ is defined on $a\in{\sf N_{I}}$, then we say that $a$ designates at $w$. If every $a\in{\sf N_{I}}$ designates at $w\in W$, then $\mathcal{I}_{w}$ is called total. We say that $\mathfrak{M}=(\mathfrak{F},\Delta,\mathcal{I})$ is a total interpretation if every $\mathcal{I}_{w}$, $w\in W$, is a total interpretation. In the sequel, we refer to partial interpretations as interpretations, and add the adjective ‘total’ explicitly whenever this is the case.
@@
The first example in Sec. 1 is thus $\mathsf{isGenChair}(\mathsf{pierre},\mathsf{kr24})$.
To avoid ambiguities, we need to use square brackets when applying negation and modal operators to assertions, as in $\lnot[C(\tau)]$ and $\Diamond_{i}[C(\tau)]$. Observe that, in an assertion of the form $\Diamond_{i}C(\tau)$, the diamond acts as a de re operator, since the concept $\Diamond_{i}C$ applies to the object, if any, designated by the term $\tau$ at the current world $w$, and the assertion is false at a world whenever $\tau$ fails to designate at $w$. On the other hand, in an expression of the form $\Diamond_{i}[C(\tau)]$, the diamond plays the role of a de dicto modality, as it refers to the world of evaluation for the whole assertion $C(\tau)$. Using the lambda abstraction notation for first-order modal logic (?), assertion $\Diamond_{i}A(a)$ corresponds to $\exists x.({\langle}\lambda y.x=y{\rangle}(a)\land\Diamond_{i}A(x))$, whereas $\Diamond_{i}[A(a)]$ stands for $\Diamond_{i}\exists x.({\langle}\lambda y.x=y{\rangle}(a)\land A(x))$; see Appendix A for details on the standard translation.
@@ 3 Reasoning Problems and Reductions
In the sequel, for the case of concept $\mathcal{C}$-satisfiability under global ontology, we will assume without loss of generality that $C$ is a concept name. Indeed, we can extend $\mathcal{O}$ with CI $A\equiv C$, for a fresh concept name $A$, and consider satisfiability of $A$ under the extended ontology, which is a model conservative extension of $\mathcal{O}$.
We begin with a few observations on polytime reductions between the concept satisfiability problems (under global ontology) for two main languages, $\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}$ and $\smash{\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}}$, and various semantic conditions, including (non-)rigid designators, total and partial interpretations, and expanding and constant domains. We also show how to eliminate the universal role using the global ontology and how to replace definite descriptions with nominals, and the other way round, which in particular means that satisfiability in the full $\smash{\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}}$ and its fragment without $\smash{\iota}$ has the same computational properties. These observations will be useful in our constructions below, where we also apply them to smaller fragments, if the results carry over. Proofs are available in Appendix B.
@@
is unsatisfiable in interpretations with the RDA, but is satisfiable otherwise, as $a$ can designate differently at different worlds. Note that an individual that fails to designate at all worlds is vacuously rigid. The following proposition shows the RDA can be enforced in interpretations by an ontology.
@@
The proof is based on the observation that CIs of the form $\{a\}\sqsubseteq\Box_{i}\{a\}$, for $i\in I$, ensure that $a$ can be made a rigid designator in any interpretation. This provides a reduction for the case of global ontology, where a given $\mathcal{O}$ is extended with such CIs for every $a$. A similar reduction is provided in Appendix B for concept satisfiability in total interpretations.
@@
We sketch the proof for the case of global ontology. Let $\mathcal{O}$ be an ontology. Consider the extension $\mathcal{O}^{\prime}$ of $\mathcal{O}$ with CIs $\top\sqsubseteq\exists u.\{a\}$, for all individual names $a$ in $\mathcal{O}$. Clearly, these CIs ensure that each $a$ designates in every accessible world. The case of concept satisfiability is shown in Appendix B.
@@
We again sketch the proof for the case of global ontology. Let $\mathcal{O}$ be an ontology. Consider $\mathcal{O}^{\prime}$ obtained from $\mathcal{O}$ by replacing every nominal $\{a\}$ in $\mathcal{O}$ with a fresh concept name $N_{a}$ and by extending the result with all CIs $N_{a}\sqsubseteq\{a\}$. It follows that every $a$ in $\mathcal{O}^{\prime}$ can designate in all worlds, but the corresponding concept $N_{a}$ may still be interpreted by the empty set in some worlds, thus reflecting the fact that $a$ in $\mathcal{O}$ could have failed to designate in those worlds. The case of concept satisfiability is dealt with in Appendix B.
@@
Next, we define normal form that will help us prove further polytime reductions, e.g., Lemma 6 and Proposition 8, and then complexity upper bounds. Let $\mathcal{O}$ be an ontology and $C$ a subconcept in $\mathcal{O}$. Denote by $\mathcal{O}[C/A]$ the result of replacing every occurrence of $C$ in $\mathcal{O}$ with a fresh concept name $A$, called the surrogate of $C$. Clearly, $\mathcal{O}[C/A]\cup\{C\equiv A\}$ is a model conservative extension of $\mathcal{O}$. We can systematically apply this procedure to obtain an ontology in normal form where connectives are applied only to concept names: e.g., definite descriptions occur only in the form of $\smash{\iota}B$, for a concept name $B$. If surrogates are introduced for innermost connectives first, then the transformation runs in polytime.
@@
We first observe that nominals can be easily encoded with definite descriptions. Indeed, given an ontology $\mathcal{O}$, take a fresh concept name $N_{a}$ for each individual name $a$ in $\mathcal{O}$, and let $\mathcal{O}^{\prime}$ be the result of replacing every occurrence of $\{a\}$ in $\mathcal{O}$ with $\{\smash{\iota}N_{a}\}$. Clearly, $\mathcal{O}^{\prime}$ is a model conservative extension of $\mathcal{O}$, and vice versa. Note that $\mathcal{O}^{\prime}$ is in the fragment $\smash{\smash{\mathcal{ML}}^{n}_{\mathcal{ALC}_{\!u}^{\smash{\iota}}}}$ without nominals, which has no distinction between partial and total interpretations, and between the RDA and no-RDA cases. Thus, we have the following result.
@@
where $a_{B}$ is a fresh individual name. Intuitively, the first CI ensures that the surrogate for $\{\smash{\iota}B\}$ belongs to $B$ and is never interpreted by more than one domain element. The second CI ensures that if $B$ is a singleton, then that element belongs to the surrogate for $\{\smash{\iota}B\}$. Formally, we show that $\mathcal{O}^{\ast}$ is a model conservative extension of $\mathcal{O}$. The case of concept satisfiability relies on normal form of concepts (Lemma 5) and is treated in Appendix B.
@@
It is known that, for the satisfiability problems, the interpretations with expanding domains can be simulated by constant domain interpretations, where a fresh concept name representing the domain is used to relativise concepts and CIs; see e.g., (?, Proposition 3.32 (ii), (iv)). We restate this standard result in our setting for completeness:
@@
In $\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}$ and $\smash{\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}}$, concept $\mathcal{C}$-satisfiability (under global ontology) with expanding domains is polytime-reducible to concept $\mathcal{C}$-satisfiability (under global ontology, respectively) with constant domain.